(4 points) Lucy factored $2,500,000 of accounts receivable with Ethel on a without recourse basis on...

  1. (4 points) Lucy factored $2,500,000 of accounts receivable with Ethel on a without recourse basis on May 1. Ethel assessed a finance charge of $15,000. Ethel retained an amount equal to 5% of the total ARs factored to cover sales returns. During May and June, customers returned merchandise to Lucy on $115,000 of credit sales. All of these returns related to receivables from the $2,500,000 pool of ARs sold. After taking the returns into consideration, Ethel collected $2,381,000 of the factored ARs. On June 30, Lucy and Ethel “settled up” meaning Lucy and Ethel paid each other any cash that was due the other.     
  1. Prepare the entry Lucy should make on May 1.
  2. Prepare the entry Lucy should make on June 30.
  3. Based on the above facts, what net profit did Ethel end up earning?
  4. Based on the above facts, what net expense did Lucy end up incurring?
Answer #1

Requirement (a)

Date General Journal Debit Credit
May 1 Cash $2,360,000
Finance Charges $15,000
Estimated sales inventory (2500000×5%) $125,000
To Accounts Receivable $2,500,000
(To record factoring of accounts receivables)

Requirement (b)

Date General Journal Debit Credit
Jun 30 Sales return $115,000
Refund Receivable $10,000
To Estimated sales returns $125,000
(To record sales retursn)
Jun 30 Cash $6,000
Bad debt expense $4,000
To Refund Receivable $10,000
(To record settlement)

Bad debt expense = Total rec factored - actual returns - Collected = 2500000-115000-2381000 = $ 4,000

Requirement (c)

Ethel earned the only income of factor charges, i.e., $15,000.

Requirement (d)

Net expense for Lucy = Factor charges + Bad debt expense + Sales returns = 15000+4000+115000 = $ 134,000.